OpenStudy (anonymous):

integral from 0 to +infinity of (xe^(x^2)) dx evaluate and tell whether it converges or diverges

OpenStudy (anonymous):

\[\int\limits_{0}^{+\infty} xe ^{-x ^{2}}dx\]

OpenStudy (anonymous):

\[\lim_{b \rightarrow +\infty} \int\limits_{0}^{b}xe ^{-x ^{2}}dx\]

myininaya (myininaya):

first let's look at \[\int\limits_{}^{}xe^{-x^2} dx\] let \[u=-x^2 => du=-2x dx\] \[\frac{-1}{2}\int\limits_{}^{} e^u du=\frac{-1}{2}e^u+C=\frac{-1}{2}e^{-x^2}+C\] so we have \[\lim_{a \rightarrow 0}\int\limits_{0}^{a}xe^{-x^2} dx=\lim_{a \rightarrow 0}[\frac{-1}{2}e^{-x^2}]_0^a\]

myininaya (myininaya):

oops infinity a-> infinity

OpenStudy (anonymous):

oh i thought i had to solve the integral by parts

myininaya (myininaya):

\[\lim_{a \rightarrow \infty}\int\limits\limits_{0}^{a}xe^{-x^2} dx=\lim_{a \rightarrow \infty}[\frac{-1}{2}e^{-x^2}]_0^a\]

myininaya (myininaya):

\[\frac{-1}{2}\lim_{a \rightarrow \infty}[e^{-a^2}-e^{-0^2}]=\frac{-1}{2}[0-e^0]=\frac{-1}{2}[0-1]\]

myininaya (myininaya):

\[=\frac{-1}{2}(-1)=\frac{1}{2}\]

OpenStudy (anonymous):

so it converges?

myininaya (myininaya):

yes we got a number so converges

OpenStudy (anonymous):

Thank you